The sum of two mubers is 36. Their difference is 12. What are the two numbers?
Found 2 solutions by dr2400, TP:
Answer by dr2400(1)

36-12=24
24-12=12
24+12=36

Answer by TP(29)

Let the two numbers be x and y.
Now their sum(total)is 36 so x+y=36 (i)
and their difference(what you get when you take one number away from the other) is 12 so x-y=12 (ii) (note: we could have said y-x=12 instead,it doesn't matter which way round you do it you will still be able to find the right answers).
So we have: x+y=36 (i)
and
x-y=12 (ii)
Now add the two equations together and we get 2x=48 (x+x=2x,+y+-y=y-y=0 and 36+12=48)
and so x=24 (a half of 48), which means that y=12 (since x+y=36).
So the two numbers are 24 and 12ANS